Lock, Statistics: Unlocking the Power of Data, 2e Chapter P, Section 1, Exercise 034 Hockey Hall of Fame From its founding through 2016, the Hockey Hall of Fame has inducted 268 players. The table below shows the number of players by place of birth and by position played. If a player is chosen at random from all player inductees nto the Hockey Hall of Fame, I t С represent the event of being born i Canada and D represent the event of being a defenseman. Hockey Hall of Fame by place of birth and position Offense Defense Goal Total 233 14 18 Canada 126 74 USA 10 Other Total 146 85 37268 Find each of the following probabilities Hockey Hall of Fame website: www.hhof.com. Correct. (a) What is the probability that an inductee chosen at random is Canadian? Round your answer to three decimal places if necessary 542 PM

Lock, Statistics: Unlocking the Power of Data, 2e (d) What is the probability that a player chosen at random is either born in Canada or a defenseman? Round your answer to three decimal places if necessary P(C or D)- the absolute tolerance is +/-0.01 E Chapter 2. Section X Incorrect. (e) what is the probability that a Canadian inductee plays defense? Round your answer to three decimal places if necessary PlD if C) = the absolute tolerance is +/-0.01 O Type here to search

P(D if C)- the absolute tolerance is +/-0.01 LINK TO TEXT 3Incorrect. (6) What is the probability that an inductee who plays defense is Canadian? Round your answer to three decimal places if necessary. e 014 P(C if D)- the absolute tolerance is +/-0.01 02 r P SHOW HINT O Type here to search
Answer #1

a)P(Canadian)=233/268=0.869

d)P(C or D) =(233+85-74)/268=0.910

e)P(D if C)=74/233=0.318

f) P(C if D)=74/85=0.871